OpenStudy (anonymous):

stephen read 90 pages in 6 days. If he continues reading at the rate, how long will it take him to read 300 pages. give an example and explain

OpenStudy (anonymous):

Same way I told you. @noterdame13 Can you try?

OpenStudy (igreen):

90 / 6 = 15 pages in 1 day 300 / 15 = ? @noterdame13

OpenStudy (anonymous):

@iGreen This would be not logical. Make her understand the question first.

OpenStudy (igreen):

This would not be logical* he*

OpenStudy (igreen):

And, it is logical..first you find out how many pages per day. 90 pages in 6 days is a ratio. 90:6 Divide that, and you get 15 pages a day. Now we can find out how many days it takes to read 300 pages by dividing 15. 300 / 15 = ? @noterdame13

OpenStudy (igreen):

*him

OpenStudy (anonymous):

first of all iam a he do you see my picture

OpenStudy (anonymous):

If Stephen reads 90 pages in 6 days,In how many days will he read 300 pages,More days or less days?, @noterdame13

OpenStudy (igreen):

Yep, I thought so. @noterdame13

OpenStudy (anonymous):

I thought It was Harry Styles in your pic. And It doesn't matter who you are. Concentrate on the Question.

OpenStudy (anonymous):

And It is Harry...

OpenStudy (anonymous):

no this is really me not harry or whatever

OpenStudy (anonymous):

LOL sure. Do you wanna learn or no?

OpenStudy (anonymous):

yea ido

OpenStudy (anonymous):

Alright,SO..

OpenStudy (anonymous):